OpenStudy (anonymous):

17x^2-11x+2=2x^2 solve the equation check your solution show work

OpenStudy (anonymous):

First subtract 2x^2 from both sides: 15x^2-11x+2=0 then use the quadratic equation a = 15 b = -11 c = 2

OpenStudy (anonymous):

17x² - 11x + 2 = 2x² 15x² - 11x + 2 = 0 15x² - 5x - 6x + 2 = 0 5x(3x - 1) - 2(3x - 1) = 0 (5x - 2)(3x - 1) = 0 x = 1/3 or x = 2/5